# Segment-to-Act: Label-Noise-Robust Action-Prompted Video Segmentation Towards Embodied Intelligence

### v1

@@ I Introduction
  • We initiate the study of noisy labels in action-based video object segmentation by formalizing text prompt noise and mask annotation noise, establishing a noisy-label taxonomy that links semantic ambiguity in language prompts with boundary imprecision in pixel masks.
  • We introduce ActiSeg-NL, the first benchmark for this setting, and unify adaptations of a broad range of noisy-label learning strategies, including Co-teaching [19], GCE [16], SCE [20], APL [21], ELR [13], and NPN [12], to the pixel-level and language-conditioned setting, providing the first large-scale comparison of noise-robust training in action-based video object segmentation.
  • Our benchmark analysis shows that robustness is not monolithic. Text noise challenges grounding and reduces foreground coverage, while Co-teaching preserves foreground regions, whereas mask noise degrades overlap, and APL improves overlap. Under mixed conditions, pixelwise losses such as GCE and SCE are more stable than sample filtering, with ELR increasing positive regions at the cost of background precision. Additionally, we propose Parallel Mask Head Mechanism (PMHM), a parallel-head consistency scheme that mitigates mask noise in action-based video object segmentation.
@@ II-B Learning with Noisy Labels
Noisy labels [26, 27], prevalent in large-scale datasets, significantly impair deep neural network performance due to their reliance on high-quality annotations. Song et al. [26] study symmetric and asymmetric noise in image classification. Luo et al. [28] leverage high-level spatial structures as supervisory signals to mitigate mislabeled annotations in segmentation tasks. Feng et al. [29] enhance weakly supervised segmentation with online pseudo-mask correction. In video object segmentation, Enki et al. [14] propose a space-mask-based framework to correct noise using spatial context, whereas Kimhi et al. [30] establish benchmarks like COCO-N and VIPER-N [30] to evaluate instance segmentation robustness against class confusion and boundary distortion. However, many of these methods operate at the image level, and most pixel-level approaches do not incorporate text guidance, so they are not directly applicable to action-based video object segmentation, which requires supervision at the pixel level conditioned on language and vision. We fill this gap by formalizing label noise for text prompts and mask annotations, as illustrated in Fig. 1, and by introducing ActiSeg-NL, the first benchmark that unifies adaptations of diverse noisy-label learning strategies for action-based video object segmentation.
@@ III Benchmark
III-A Text Prompt Noise
Construction. We generate text prompt noise on the VISOR training split with two components. First, we construct symmetric noise by randomly flipping referent objects in text prompts, such as replacing “container” with “fridge” in the prompt “celery used in the action of put celery back into fridge” with probabilities of $20\%$, $40\%$, or $60\%$, which represent low, medium, and high noise levels commonly used in noisy label learning to assess robustness, following [21, 13]. The target category is randomly selected from a predefined set to ensure equal replacement probability. Second, we substitute nouns with synonymous or related terms within the same category, such as replacing “container” with “food container” or “cheese container” to increase semantic complexity, by uniformly sampling the replacement from $\mathcal{M}(\text{class}_{\mathcal{C}})$ when the class flipping condition is not met, as described in Alg. 1. The whole procedure is summarized in Alg. 1.
@@ III-A Text Prompt Noise
Analysis. We diagnose distribution shifts induced by text prompt noise. To validate this noise strategy, we analyze category distributions in the original and noisy training sets. Fig. 2 compares major categories (proportion $1\%$ in the original set) and a summarized “Others” category across clean data and $20\%$, $40\%$, and $60\%$ noise rates. The statistics shows that “left hand” and “right hand” proportions decrease from $14.58\%$ and $15.60\%$ to $6.08\%$ and $6.39\%$ at $60\%$ noise rate, respectively, while “Others” increases from $44.43\%$ to $75.47\%$. This shift validates the disruptive effect of our noise strategy, as depicted in Fig. 1, establishing ActiSeg-NL as a dedicated benchmark for exploring text prompt noise in video segmentation.
@@ III-B Mask Annotation Noise
Analysis. By merging the dilated per-object masks with a first hit rule under a fixed object order, overlaps create boundary competition that resembles the blur observed in crowdsourced annotations, yet the resulting noise is more challenging than typical crowdsourced boundary errors and grows stronger as the dilation kernel increases. The severity of the generated mask noise is directly controlled by the dilation kernel size. As illustrated in the right of Fig. 3, increasing the kernel size from $9$ to $21$ monotonically degrades annotation quality.
@@ IV Methodology
Setup and Notation We study action-based video object segmentation [6] with noisy supervision. The task is to generate a per-frame binary mask identifying the active object, given a video clip and a textual prompt. Formally, for each pixel $(x,y)$ in a frame $t$, the model predicts a foreground probability $p_{t}(x,y)\in[0,1]$. This prediction is trained against a corresponding ground truth label $y_{t}(x,y)\in\{0,1\}$, which may be noisy. Training minimizes a frame-averaged pixel-wise objective to ensure reliable localization for embodied manipulation.
@@ IV-A Adapted Robust Learners for Pixel-level and Language-conditioned Segmentation
Early Learning Regularization (ELR) [13] is designed to prevent the model from memorizing noisy labels by regularizing its predictions to be consistent with their historical average. It combines a primary classification loss with a temporal agreement regularizer. The primary loss is the focal loss [31], which uses weights $w^{+}=\alpha(1-p_{t}^{(n)})^{\gamma}$ and $w^{-}=(1-\alpha)(p_{t}^{(n)})^{\gamma}$ to down-weight well-classified examples:
@@
NPN [12] integrates partial label learning, negative learning, and dual-view consistency at the pixel level. For each pixel form candidate indicators $A_{t}^{(1,n)}=y_{t}+\mathbb{I}\!\bigl[p_{t}^{(n)}>0.5\bigr]$ and $A_{t}^{(0,n)}=(1-y_{t})+\mathbb{I}\!\bigl[p_{t}^{(n)}\leq 0.5\bigr]$ and accumulate counts $M_{t}^{(1)}\leftarrow M_{t}^{(1)}+A_{t}^{(1,n)}$ and $M_{t}^{(0)}\leftarrow M_{t}^{(0)}+A_{t}^{(0,n)}$. The proxy label is $y_{t}^{\star}=\mathbb{I}[M_{t}^{(1)}\geq M_{t}^{(0)}]$ and the reliability is $w_{t}=\dfrac{\max(M_{t}^{(1)},M_{t}^{(0)})}{M_{t}^{(1)}+M_{t}^{(0)}}$. The partial label loss $\ell_{\text{PLL}}^{\,txy}$ is
@@
the negative learning loss $\ell_{\text{NL}}^{\,txy}$ is
@@
$-\bigl(1-\mathbb{I}[A_{t}^{(0,n)}\geq 1]\bigr)\log p_{t}^{(n)},$
and the dual-view consistency loss $\ell_{\text{CR}}^{\,txy}$ uses a weak-to-strong cross-entropy with hard targets from the weak view. The final loss averages $\ell_{\text{PLL}}^{\,txy}+\alpha\,\ell_{\text{NL}}^{\,txy}+\beta\,\ell_{\text{CR}}^{\,txy}$ over pixels and frames.
IV-B Proposed PMHM Consistency with Main and Auxiliary Mask Heads
@@
The total segmentation loss $L_{\text{seg}}$ combines a hard mask loss $L_{\text{hard}}$ [6] on confident pixels (those outside $U_{t}$) with our consistency terms:
@@ V-D Effects of Mask Annotation Noise
With $k{\times}k$ morphological dilation applied to ground-truth masks (Table II), performance degrades chiefly on positive-region metrics and the aggregate score: p-mIoU drops steeply at moderate noise (65.065.055.755.7 at $9{\times}9$) and then continues to fall with diminishing increments ($49.1$ at $15{\times}15$, $44.8$ at $21{\times}21$), while gIoU declines from 70.770.768.268.263.763.760.360.3. Negative-region metrics (↓ better) improve at moderate noise but show a non-monotone rebound at the highest level: n-mIoU 19.719.713.713.713.313.313.913.9 and n-cIoU 34.134.120.320.319.719.729.929.9. This pattern indicates that boundary dilation first discourages spurious background activations, pushing the model toward conservative masks with fewer false positives. Yet, once the noisy boundary band becomes too wide, supervision near edges turns conflicting, inducing boundary drift and shape jitter that hurt positive overlap and increase large area background overlap (hence the n-cIoU rebound). Acc changes little (81.781.779.279.2), suggesting the classifier is relatively insensitive compared with segmentation overlap metrics. At matched severities, mask noise is also more damaging than text noise (e.g., gIoU $60.3$ at $21{\times}21$ vs. $67.1$ at Name-60; cf. Table I), underscoring the need for boundary-focused robustness and for reporting p/n-partitioned metrics rather than relying on a single aggregate score.
@@ V-E Evaluation of Robustness Strategies under Combined Noise Conditions
Observation 2 is a gradient-conflict effect under mixed noise. Prompt mis-grounding encourages conservative masks (shrinking foreground to avoid false activations), whereas boundary dilation encourages expansion near edges; pixel-wise robust losses (GCE/SCE/APL) and dual-view consistency (NPN) better reconcile these opposing update directions than sample filtering alone. This interpretation is consistent with the design goals of GCE/SCE/APL [16, 20, 21] and NPN [12], now evidenced in a multimodal, pixel-level setting.
Observation 3 concerns metric sensitivity and benchmark coverage. Because positive and negative regions move in opposite directions under mixed noise, aggregate gIoU can mask failures; reporting p-/n-partitioned mIoU/cIoU surfaces boundary-specific errors and background precision. The consistent re-ranking between Table III and Table IV validates the need for our multimodal protocols and partitioned metrics to fully characterize robustness in action-based video object segmentation.
@@ V-F PMHM under Boundary and Mixed Noise
As PMHM is designed to address corrupted mask labels, we evaluated its performance under boundary and mixed noise conditions. In the boundary-only noise setting (kernel sizes $9$, $15$, and $21$), PMHM consistently reduces background errors and improves global scores over the baseline. At the highest noise level ($21{\times}21$), it achieves a gIoU of $61.7$, outperforming the baseline’s $60.3$ (Table II). In the mixed-noise setting (Name-60 + Mask-21), where severe text noise is introduced, PMHM’s performance (gIoU $59.9$) does not surpass the baseline FLac (gIoU $60.5$), as shown in Table IV. This confirms that while PMHM enhances robustness against its primary target, boundary corruption, its benefits are less pronounced when combined with significant text prompt noise.

### v2

@@ I Introduction
  • We initiate the study of noisy labels in action-based video object segmentation by formalizing text prompt noise, mask annotation noise, and mixed noise, establishing a noisy-label taxonomy that links semantic ambiguity in language prompts with boundary imprecision in pixel masks.
  • We introduce ActiSeg-NL, the first benchmark for this setting, and unify adaptations of a broad range of diverse noisy-label learning strategies, including Co-teaching [10], GCE [36], SCE [28], APL [34], ELR [18], and NPN [26], to the pixel-level and language-conditioned setting, providing the first large-scale comparison of noise-robust training in action-based video object segmentation.
  • Our benchmark analysis shows that robustness is not monolithic. Text noise challenges grounding and reduces foreground coverage, while Co-teaching preserves foreground regions, whereas mask noise degrades overlap, and APL improves overlap. Under mixed conditions, pixelwise losses (e.g., GCE, SCE) are more stable than sample filtering, with ELR increasing positive regions at the cost of background precision. Additionally, we propose Parallel Mask Head Mechanism (PMHM), a parallel-head consistency scheme that mitigates mask noise in action-based video object segmentation.
@@ II-B Learning with Noisy Labels
Noisy labels [27, 17], prevalent in large-scale datasets, significantly impair deep neural network performance due to their reliance on high-quality annotations. Song et al. [27] study symmetric and asymmetric noise in image classification. Luo et al. [20] leverage high-level spatial structures as supervisory signals to mitigate mislabeled annotations in segmentation tasks. Feng et al. [8] enhance weakly supervised segmentation with online pseudo-mask correction. In video object segmentation, Enki et al. [4] propose a space-mask-based framework to correct noise using spatial context, whereas Kimhi et al. [13] establish benchmarks like COCO-N and VIPER-N [13] to evaluate instance segmentation robustness against class confusion and boundary distortion. However, many of these methods [27, 17] operate at the image level, and most pixel-level approaches [32, 8, 4] do not incorporate text guidance, so they are not directly applicable to action-based video object segmentation, which requires supervision at the pixel level conditioned on language and vision. We fill this gap by formalizing label noise for text prompts and mask annotations, as illustrated in Fig. 1, and by introducing ActiSeg-NL, the first benchmark that unifies adaptations of diverse noisy-label learning strategies for action-based video object segmentation.
@@ III Benchmark
III-A Text Prompt Noise
Construction. We generate text prompt noise for the VISOR training split on target objects with two components. First, symmetric noise is built by randomly flipping referent objects in text prompts, such as replacing “container” with “fridge” in the prompt “container used in the action of take celery” with probabilities of $20\%$, $40\%$, or $60\%$, which represent low, medium, and high noise levels commonly used in noisy label learning to assess robustness, following [34, 18]. The target category is randomly selected from a predefined set to ensure equal replacement probability. Second, we substitute nouns with synonymous or related terms within the same category, such as replacing “container” with “food container” or “cheese container” to increase semantic complexity, by uniformly sampling the replacement from $\mathcal{M}(\text{class}_{\mathcal{C}})$ when the class flipping condition is not met, as described in Alg. 1. The whole procedure is summarized in Alg. 1.
Analysis. We diagnose distribution shifts induced by text prompt noise. To validate this noise strategy, we analyze category distributions in the original and noisy training sets. Fig. 2 compares major categories (proportion $1\%$ in the original set) and a summarized “Others” category across clean data and $20\%$, $40\%$, and $60\%$ noise rates. These statistics show that “left hand” and “right hand” proportions decrease from $14.58\%$ and $15.60\%$ to $6.08\%$ and $6.39\%$ at the $60\%$ noise rate, respectively, while “Others” increases from $44.43\%$ to $75.47\%$. This shift validates the disruptive effect of our noise strategy, as depicted in Fig. 1, establishing ActiSeg-NL as a dedicated benchmark for exploring text prompt noise in video segmentation.
@@ III-B Mask Annotation Noise
Analysis. By merging the dilated per-object masks with a first hit rule under a fixed object order, overlaps create boundary competition that resembles the blur observed in crowdsourced annotations, yet the resulting noise is more challenging than typical crowdsourced boundary errors and grows stronger as the dilation kernel increases. The severity of the generated mask noise is directly controlled by the dilation kernel size. Increasing the kernel size from $9$ to $21$ monotonically degrades annotation quality (Fig. 3, right).
@@ IV Methodology
Setup and Notation. We study action-based video object segmentation [24] with noisy supervision. The task is to generate a per-frame binary mask identifying the active object, given a video clip and a textual prompt. Formally, for each pixel $(x,y)$ in a frame $t$, the model predicts a foreground probability $p_{t}(x,y)\in[0,1]$. This prediction is trained against a corresponding ground truth label $y_{t}(x,y)\in\{0,1\}$, which may be noisy. Training minimizes a frame-averaged pixel-wise objective to ensure reliable localization for embodied manipulation.
@@ IV-A Adapted Robust Learners for Pixel-level and Language-conditioned Segmentation
ELR [18] prevents noisy label memorization by regularizing its predictions toward their historical average. It combines a primary classification loss with a temporal agreement regularizer. The primary loss is the focal loss [15], which uses weights $w^{+}{=}\alpha(1{-}p_{t}^{(n)})^{\gamma}$ and $w^{-}{=}(1{-}\alpha)(p_{t}^{(n)})^{\gamma}$ to down-weight well-classified examples:
@@
NPN [26] integrates partial label learning, negative learning, and dual-view consistency at the pixel level. For each pixel, we form candidate indicators $A_{t}^{(1,n)}=y_{t}+\mathbb{I}\!\bigl[p_{t}^{(n)}>0.5\bigr]$ and $A_{t}^{(0,n)}=(1-y_{t})+\mathbb{I}\!\bigl[p_{t}^{(n)}\leq 0.5\bigr]$ and accumulate counts $M_{t}^{(1)}\leftarrow M_{t}^{(1)}+A_{t}^{(1,n)}$ and $M_{t}^{(0)}\leftarrow M_{t}^{(0)}+A_{t}^{(0,n)}$. The proxy label is $y_{t}^{\star}=\mathbb{I}[M_{t}^{(1)}\geq M_{t}^{(0)}]$ and the reliability is $w_{t}=\dfrac{\max(M_{t}^{(1)},M_{t}^{(0)})}{M_{t}^{(1)}+M_{t}^{(0)}}$. The partial label loss $\ell_{\text{PLL}}^{\,txy}$ is
@@
and the negative learning loss $\ell_{\text{NL}}^{\,txy}$ is
@@
$-\bigl(1-\mathbb{I}[A_{t}^{(0,n)}\geq 1]\bigr)\log p_{t}^{(n)}.$
Dual-view consistency $\ell_{\text{CR}}^{\,txy}$ applies weak-to-strong cross-entropy with hard targets from the weak view. The final loss averages $\ell_{\text{PLL}}^{\,txy}+\alpha\,\ell_{\text{NL}}^{\,txy}+\beta\,\ell_{\text{CR}}^{\,txy}$ over pixels and frames.
@@
IV-B Proposed PMHM Consistency with Main and Auxiliary Mask Heads
@@
The loss $L_{\text{seg}}$ combines a hard mask loss $L_{\text{hard}}$ [24] on confident pixels (outside $U_{t}$) with our consistency terms:
@@ V-D Effects of Mask Annotation Noise
With $k{\times}k$ morphological dilation applied to ground-truth masks (Table II), performance degrades chiefly on positive-region metrics and the aggregate score: p-mIoU drops steeply at moderate noise (65.065.055.755.7 at $9{\times}9$) and then continues to fall with diminishing increments ($49.1$ at $15{\times}15$, $44.8$ at $21{\times}21$), while gIoU declines from 70.770.768.268.263.763.760.360.3. Negative-region metrics (↓ better) improve at moderate noise but show a non-monotone rebound at the highest level: n-mIoU 19.719.713.713.713.313.313.913.9 and n-cIoU 34.134.120.320.319.719.729.929.9. This pattern indicates that boundary dilation first discourages spurious background activations, pushing the model toward conservative masks with fewer false positives. Yet, once the noisy boundary band becomes too wide, supervision near edges turns conflicting, inducing boundary drift and shape jitter that hurt positive overlap and increase large area background overlap (hence the n-cIoU rebound). Acc changes little (81.781.779.279.2), suggesting the classifier is relatively insensitive compared with segmentation overlap metrics. At matched severities, mask noise is also more damaging than text noise (e.g., gIoU $60.3$ at $21{\times}21$ vs. $67.1$ at Name-60; cf. Table I), underscoring the need for boundary-focused robustness and for reporting p/n-partitioned metrics over a single aggregate score.
@@ V-E Evaluation of Robustness Strategies under Combined Noise Conditions
Observation 2 is a gradient-conflict effect under mixed noise. Prompt mis-grounding encourages conservative masks (shrinking foreground to avoid false activations), whereas boundary dilation encourages expansion near edges; pixel-wise robust losses (GCE [36], SCE [28], APL [34]) and dual-view consistency (NPN [26]) better reconcile these opposing update directions than sample filtering alone. This interpretation matches the design goals of GCE/SCE/APL and NPN, now evidenced in a multimodal, pixel-level setting.
Observation 3 concerns metric sensitivity and benchmark coverage. Since positive and negative regions shift oppositely under mixed noise, aggregate gIoU can mask failures; reporting p-/n-partitioned mIoU/cIoU surfaces boundary-specific errors and background precision. The consistent re-ranking between Table III and Table IV validates the need for our multimodal protocols and partitioned metrics to fully characterize robustness in action-based video object segmentation.
@@ V-F PMHM under Boundary and Mixed Noise
As PMHM is designed to address corrupted mask labels, we evaluated its performance under boundary and mixed noise conditions. In the boundary-only noise setting (kernel sizes $9$, $15$, and $21$), PMHM consistently reduces background errors and improves global scores over the baseline. At the highest noise level ($21{\times}21$), it achieves a gIoU of $61.7$, outperforming the baseline’s $60.3$ (Table II). In the mixed-noise setting (Name-60 + Mask-21), where severe text noise is introduced, PMHM’s performance (gIoU $59.9$) does not surpass the baseline FLac (gIoU $60.5$), as shown in Table IV. This confirms that while PMHM resists boundary corruption, its gains fade when combined with significant text noise.